What is the total electric flux through a closed surface containing a 2.0 μC charge?

1
Expert's answer
2022-02-16T08:34:53-0500

Given:

"q=2.0*10^{-6}\:\rm C"

"\epsilon_0=8.85*10^{-12}\:\rm F/m"


The Gauss's law says

"\phi=\oint {\bf E}d{\bf A}=\frac{q}{\epsilon_0}"

Hence

"\phi=\frac{2.0*10^{-6}\:\rm C}{8.85*10^{-12}\:\rm F/m}=2.3*10^5\: \rm V\cdot m"